This is a discrete math question:

Exercise 5. Let A and B be sets. Prove or disprove: AAB| = |A - B+B - AL. Claim. Proof
Answer #1

Any query in this then comment below.. i will explain you..

Bove - I AAB) - LA-B) & TB-A) Prosty A ABE (AB) (BA) مانام A-B = A - CAAB) B-A= B- (AMB) (A-B) n (BA) = d so . & AAB= (A-B) (